OpenStudy (anonymous):

Will fan and medal What is the value of 2(4i2) + 22? -14 -10 14 30 54

OpenStudy (anonymous):

@TheEdwardsFamily

OpenStudy (anonymous):

@annah.bee

OpenStudy (anonymous):

@annah.bee can you please help

OpenStudy (annah.bee):

is this on k12? cuz if so i can pull it up

OpenStudy (anonymous):

not really it's NWEA

OpenStudy (anonymous):

i^2? or 4i*2?

OpenStudy (annah.bee):

sorry idk what the answer is :(

OpenStudy (anonymous):

i^2

OpenStudy (anonymous):

14

OpenStudy (anonymous):

How do you get 14 can you explain

OpenStudy (anonymous):

4i^2=-4 -4*2=-8 -8+22=14

OpenStudy (anonymous):

ohhhhh I understand now Thank you so much

OpenStudy (anonymous):

Sure